OpenStudy (anonymous):

It costs 22.00 per hour plus a flat fee of 33.00 for a plumber to make a house call. After writing an equation for his situation, suppose the total to have a plumber come to a house is 231. How many hours did the plumber work?

OpenStudy (campbell_st):

well let C = cost and H = hours.. the plumber gets $33 for just turning up and 22 per hour so C = 22h + 33

OpenStudy (anonymous):

How many hours did he work? I got 7 cause I divided 231/33

OpenStudy (anonymous):

The answers given are 6hr, 19hr, 16 hr or 9 hr.

OpenStudy (campbell_st):

well you have an equation to solve \[132 = 22h + 33\] 1st step is to subtract 33 from both sides of the equation... \[99 = 22h\] now divide both sides of the equation by 22 the answer is the number of hours

OpenStudy (anonymous):

How did you get 132? Do you suppose to use 231?

OpenStudy (campbell_st):

oops typo so 231 = 22h + 33 subtract 33 from both sides 198 = 22h divide both sides of the equation by 22 this is the number of hours...

OpenStudy (anonymous):

I got 9 hours

OpenStudy (campbell_st):

well that makes sense..

OpenStudy (anonymous):

c=22h+33 231=22h+33 231-33=22h 198=22h h=198/22=9hours